A rolled steel Joist of I-Section has flange length of 300 mm. wide and 20 mm thick with a web thickness of 20 mm. and overall depth of I-Section is 600 mm. If this beam carries a UDL of 40 KN/m over the simply supported beam of span 10 m, find the maximum stress produced in the beam.                                                                                             


1
Expert's answer
2021-12-03T04:48:02-0500

From bending equation

"\frac{M}{I}=\frac{\sigma}{y}=\frac{E}{R}"

"\bar{y}=\frac{\sum{y_iA_i}}{\sum{A_i}}=\frac{6960000}{23200}=300mm"

"\sum F_y=0"

"R_A+R_B=400"

"\sum{M_A}=0"

"10R_B=2000\implies R_B=200"

RA=400-200=200

Maximum moment occurs at C

"{M_C}=200*5-40*5*\frac{5}{2}"

MC=500

"I_{xx}=1.303*10^9"

"\frac{M_{max}}{I}=\frac{\sigma_{max}}{\bar{y}}"

"\sigma_{max}=\frac{M_{max}*\bar{y}}{I}"

="\frac{500*10^3*10^3*300}{1.303*10^9}"

=115.118N/mm2